AmeenaSarwary:

Which of the binomials below is a factor of this expression? 25x2 - 4y2 z2 A. 5x - 2yz B. 5x - 4yz C. 5x2 - 4y2 z2 D. 5x2 - 2y2 z2c

Nnesha:

Difference of squares \[\huge\rm a^2 - b^2 = (a-b)(a+b)\]

Nnesha:

\[\large\rm a^2-b^2 =(\sqrt{a^2}\color{red}{-}\sqrt{b^2})(\sqrt{a^2}\color{blue}{+} \sqrt{b^2})= (a\color{red}{-}b)(a\color{red}{+}b)\]